OpenStudy (anonymous):

Which of the ordered pairs, (4,3), (7,5), and (4,14) are solutions for the equation 3x + y = 26? A. (7,5) B. (4,3), (4,14) C. (7,5), (4,3) D. (7,5), (4,14)

OpenStudy (lgbasallote):

substitute each one for (4,3) 3(4) + 3 = 26? 12 + 3 is not equal to 26 for (7,5) 3(7) + 5 = 26? 21 + 5 = 26 so )7,5) is your o.p.

OpenStudy (anonymous):

@vince33. given the vertices substitute it to the equation, use trial and error method. .

OpenStudy (anonymous):

thank you